OpenStudy (anonymous):

exponential function - F(x)=3^x-2, what is x when Y=1/3

OpenStudy (anonymous):

not really sure the method behind it, but when y=1/3 your equation looks like 1/3=3^x-2 1/3 is equal to 3^-1 so you can set up an equation with just the exponents -1=x-2 then solve for x x=1

OpenStudy (anonymous):

actually now that i think about it, that is the method to use as well, try to set up bases of the exponents that are the same, so you can then just focus on the exponents in equation form